Status Report

Hubble Space Telescope Daily Report #2760 12/01/00

By SpaceRef Editor
December 1, 2000
Filed under ,

PERIOD COVERED:  0000Z (UTC) 11/30/00 – 0000Z (UTC) 12/01/00
Daily Status Report as of 336/0000Z
    1.1 Completed Three Sets of WF/PC-2 8826 (Cycle 9 Supplemental Darks pt 1)
        The WF/PC-2 was used to perform a T dark calibration program that obtains three dark frames every day in order to provide data for monitoring and characterizing the evolution of hot pixels.  There were no reported anomalies.
    1.2 Completed STIS/CCD 8661 (UV Spectroscopy of the Giant Planet Atmospheres)
        The Space Telescope Imaging Spectrograph (CCD) was used to observe the atmosphere of Uranus in the ~1600-3100 Angstrom spectral
region.  Later, the remaining giant planets, Jupiter, Saturn, and Neptune will be observed.  Analyses of these observations will be used to determine the abundance and altitude profiles of gaseous absorbers and aerosols from the millibar region to hundreds of millibars.  The proposal completed with no reported problems.
    1.3 Completed Three Sets of STIS/CCD 8837 (CCD Dark Monitor-Part 1)
        The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks.  The proposal completed with no reported problems.
    1.4 Completed WF/PC-2 8683 (Imaging Of Brightest Cluster Galaxies: The High End Of The Black Hole Mass Distribution)
        The WF/PC-2 was used to make kinematic black hole detections in galaxies to decide whether they indicate that the mass correlates with both optical luminosity and radio power.  The observation completed with no reported problems.
    1.5 Completed STIS/CCD 8562 (Probing the Large Scale Structure: Cosmic Shear Observations)
        The Space Telescope Imaging Spectrograph (CCD) was used to probe the distortion of light bundles from distant galaxies, looking at the statistical properties of the intervening inhomogeneous {dark} matter distribution.  The proposal completed nominally.
    1.6 Completed WF/PC-2 8389 (Illuminating the Galactic Dark Matter)
        The WF/PC-2 was used to re-image the HDF in F606W in order to confirm and better determine the proper motions of four MACHO candidates discovered from the comparison of the original December 1995 HDF frames with WFPC2 F814W frames exposed in December 1997.  The observations completed with no reported problems.
    1.7 Completed Four Sets of STIS/CCD 8260 (Searching for the Hydrogen Reionization Edge of the Universe at 5HST)
        The Space Telescope Imaging Spectrograph (CCD) was used to observe one of 36 parallel orbits {4–5 fields of 5–8 orbits each} to constrain the H Lyman-edge in emission that marks the transition from a neutral to a fully ionized IGM at a predicted zion~eq5–15.  This edge is due to recombination from the H Lyman series and Lyman continuum, and can be used to constrain zion, one of the most important unknown quantities in large scale structure and cosmology.  The proposal completed nominally.
    1.8 Completed STIS/CCD 8838 (Bias Monitor – Part 1)
        The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1×1, 1×2, 2×1, and 2×2 bin settings at gain=1, and 1×1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns.  The proposal completed nominally.
    1.9 Completed Five Sets of WF/PC-2 8699 (The Origin of Short-Period Comets)
        The WF/PC-2 was used to detect and characterize cometary nuclei in order to determine the basic physical properties of a large fraction of the population of short-period comets.  By acquiring statistically significant data, we can study the origin of this family of comets and test the hypothesis that they are collisional fragments from the Kuiper Belt Objects.  As described in HSTAR 7982 and 2.1, the acquisition for the final iteration of this proposal failed, all observations were lost.  Otherwise, the observations completed with other no reported problems.
    1.10 Completed Two Sets of STIS/MA2 8843 (Cycle 9 MAMA Dark Measurements)
        The Space Telescope Imaging Spectrograph (MA2) was used to perform the routine monitoring of the MAMA detector dark noise.  The proposal completed nominally.
    2.1 Guide Star Acquisitions:
      Scheduled Acquisitions: 11
Successful: 10
        Per HSTAR 7982, the acquisition at 335/214034Z failed when the search radius limit was exceeded for FGS-2.  A following map showed errors of 6.747, -2.719, and -7.079.  The proposal detailed in 1.9 was affected.   Scheduled Re-acquisitions:  4
Successful:  4
    2.2 FHST Updates:
Scheduled: 21
Successful: 21
    2.3 Operations Notes:
        HST operations was conducted from the Science Institute as GSFC Internal Simulation #2 (EVA-1) was conducted for much of the day from the GSFC STOCC.
        Using ROP SR-1A, the SSR EDAC error counter was cleared four times.
        HSTAR 7981 documents that at 335/022153Z the coarse sun sensor #1 temperature (GCSS1T) flagged our-of-limits low (-45.0897) for one telemetry sample.
        The engineering status buffer limits were adjusted at 335/2153Z, per ROP DF-18A.
        SSA transmitter #2 was turned on at 336/0435Z and turned off at 336/0445Z with ROP IC-2.
        CCS string "B" (v3.3) was configured for prime operations at 336/0303Z.  When there was a core data server problem, CCS string "A" (v3.2) was configured for prime operations.  Later, at 336/0400Z, CCS string "B" was configured for prime operations.
        The STIS MCE-1 reset at 336/080327Z while the low voltage was on and while inside an SAA interval.  The STIS flight software error counter was reset at 336/0907Z as directed by ROP NS-12.  MAMA-1 will be recovered at 336/1342Z via normal SMS commanding.
        After analysis, the on-board automatic gyro bias computations were enabled at 336/1137Z.
        Continuation of normal science observations and calibrations.

SpaceRef staff editor.